OpenStudy (anonymous):

What is the solution to the rational equation x over x squared minus 9 minus 1 over x plus 3 equals 1 over 4x minus 12 ? x = -9 over 7 x = 15 x = 15 over 7 x = 9

OpenStudy (anonymous):

\[\frac{ x }{ x ^{2}-9 }-\frac{ 1 }{ x+3 }=\frac{ 1 }{ 4x-12 }\] \[\frac{ x }{ (x-3)(x+3) }-\frac{ 1 }{ (x-3) }=\frac{ 1 }{ 4(x-3) }\] multiply both sides by (x-3) \[\frac{x}{(x+3)}-1=\frac{1}{4}\] add 1 to both sides \[\frac{x}{x+3}=1\frac{1}{4}\] multiply both sides by (x+3) \[x=3\frac{3}{4}+1\frac{1}{4}x\] subtract \[1\frac{1}{4}x\] from both sides \[-\frac{1}{4}x=3\frac{3}{4}\] divide both sides by \[-\frac{1}{4}\] x=-15 ik its not one of the choices but that is what I got